ied Exercise 9-6 rales Company acqures a delivery truck at a cost of $38,000. The truck is expected to have a savage value of $6,000 at the end of ts 4-year unelife. Assuming the decining baiance depreciation rate is double preciation for the frst and second years under the dedining-balance method ick if you would like to Show Woerk for this question: Qeenhon Wack
Answer #1
Depreciation rate = 1/4*2 = 50%
Annual depreciation expense:
Year 1 19000 =38000*50%
Year 2 9500 =(38000-19000)*50%